Instrument Log

60 M ECOR   C1               CF Lamont, OK

Date

Operator

Comments

7/12/01

JEM

Changed Bernoulli Disk.  RTS @2000 GMT.

6/28/01

JEM

Changed Bernoulli Disk.  Unable to establish communications with the 60M Ecor because the Bernoulli Drive will not spin down.  Several attempts to power cycle the DCU were unsuccessful.  R/R *Bernoulli Drive.  RTS @18   GMT.   *Old – 7X1G5C    *New – 7X02TR

5/17/01

MK

PM and changed out 60M disk. RTS at 1915 GMT.

4/20/01

ZW

PM and disk change out. The 60M ECOR sonic would not respond to commands from the laptop computer. Power cycled the sonic at 1430 GMT. Communications were restored. The 60M Hygrometer also is working fine, however there are no instruments on the 25 M boom at this time. RTS 1500 GMT.

4/5/01

MK

PM. Changed out Bernoulli disk. The Hygrometer sensor is not working at this time. RTS at 1945 GMT.

3/22/01

MK

PM and changed out Bernoulli disk, the HYGROMETER is not working at this time. RTS at 1855 GMT.

3/8/01

JEM

Changed Bernoulli Disk. RTS @2015 GMT.

2/1/2001

MR/KT

Lowered the SW (B) side of the 60M tower from 2140 GMT to 2125 GMT to secure a loose hygrometer cable at the 60M level. All instruments on the SW side of the tower were affected during this time.

12/28/00

ZW

PM and disk change out. There are no instruments on the 25M level. The 60M ECOR would not respond to commands from the laptop computer. Power cycled the 60M ECOR system at 1930. Communications were restored. Status OK. RTS 1945 GMT.

8/11/00

ZW

PM and disk change out. Status OK. RTS 1445 GMT.

6/23/00

ZW

The 60M ECOR was missing it's DCU. The C1 ECOR *DCU was picked up at storage, & installed at 2000 GMT. The two disks for the 25M level, & 60M level both stopped spinning. The new DCU communicated with the laptop however there are only instruments on the 60M level. The 60M sonic, & hygrometer both had good readings. There are no spare C1 sonics in spares at this time. When a spare C1 sonic becomes available it, and a hygrometer will be installed at the 25M level. RTS 2100 GMT. *New - 940101

2/25/2000

ZW

Attempted to monitor the 60M ECOR DCU at 1930 GMT, however the DCU would not communicate with the laptop computer. Several attempts to power cycle the DCU were not successful. Removed the 60M *DCU at 2000 GMT. Do to a lack of a spare DCU none was installed at this time. The 60M ECOR is down at this time. RTS 2045 GMT. *Old - 940101

8/13/99

ZW

PM and disk change out. The 60 SONIC was operating in good order. The 60M Hygrometer would not respond to operator commands from the laptop computer. Power cycled the 60M Hygrometer at 1345 GMT. Communications were restored. RTS 1400 GMT.

7/16/99

MK

The 60 M Hygrometer would not respond to operator commands. Lowered tower at 1430 GMT and R/ R *Hygrometer, communications were restored and all readings were normal. *Old - 940213. *New - 960202.

7/2/99

ZW

The C1 60/25 meter ECOR has been down because there was no C1 DCU in spare. The C1 ECOR *DCU came in so Site Ops took it up to the tower, and installed it at 1720. Site Ops also installed a **Hygrometer on the 60 meter platform. The 60M and 25M meter SONIC's worked fine, however the 25M and 60M Hygrometers would not communicate with the laptop computer. Site Ops power cycled the system several times with no success. RTS 2020 GMT. *New - 960201 **New - 940213

5/19/99

5/20/99

CAM/DH

Dick Hart and Gary Zimmerman of ATI on site to T/S persistent 25/60M ECOR problems. The hygrometer appeared to be working O.K. locally but the sonic sensor is dropping characters when sending data to the DCU. The data as collected by the DCU has been bad ever since system installation. Lowered the tower booms at 1755 GMT on 5/19/99. Could not establish communications with the hygrometer at ground level. The sonic sensor continues to intermittently drop characters from the data stream and the data collected by the DCU is unusable. Gary determined that the data problems are most likely due to the DCU software. Removed the *DCU and will return to ATI for diagnostics and software corrections. Attempted to install the spare DCU but encountered similar problems. Spare DCU will also be returned to ATI. Removed the **hygrometer to return to ATI for repair. Raised booms at 1755 GMT on 5/20/99. The sonic sensor remains in place but the ECOR is off line at this time. *Old - 940101 **Old - 940214

4/30/99

MK

The Hygrometer is not responding to operator commands. Lowered tower at 1515 GMT and R/R *Hygrometer. Raised tower at 1730 GMT. RTS 1730 GMT. *Old - 960202 *New - 940214

3/12/99

MK

PM and changed out bernoulli disk. The SONIC status is normal. The HYGROMETER is not responding to operator commands. P/C HYGROMETER at 1515 GMT, still no change. Unable to replace due to lack of spares. RTS at 1535 GMT.

2/12/99

MK

PM and changed out bernoulli disk. The 60M HYGROMETER is not responding to operator commands. Power cycled several times at 1715 GMT, still no response. Lowered tower at 1730 GMT and R/R *HYGROMETER. Raised tower at 1815 GMT. At this time, the 60M SONIC and HYGROMETER are operating properly. RTS at 1830 GMT. *Old - 940214 *New - 960202

1/29/99

ZW

PM and disk change out. The 25M SONIC, and Hygrometer are missing. Due to lack of spare sensor no corrective action taken at this time. 60M SONIC, Hygrometer are working in fine order. Status ok. RTS 1515 GMT.

1/15/99

ZW

PM and disk change out. The 60M SONIC and 60M Hygrometer are working fine. The 25M SONIC, and 25M Hygrometer are down at this time and will be down until spares become available. RTS 1445 GMT.

12/30/98

ZW

PM and disk change out. The 25M SONIC, and 60M Hygrometer would not respond to operator commands from the laptop computer. The 25M Hygrometer is missing at this time. Power cycled the 25M SONIC, and the 60M Hygrometer several times however this did not fix this problem. Due to lack spare sensors no more corrective action was taken at this time. The 60M SONIC is the only sensor working at this time. RTS 2000 GMT.

12/17/98

ZW

PM & disk change out. The C1 ECOR 1a SONIC, and 1c Hygrometer would not respond to operator commands. Power cycled the 1a SONIC, and 1c Hygrometer at 1940 GMT. This did not fix the problem. Because of no spare C1 SONIC there was no more corrective action taken at this time. RTS 1955 GMT

12/7/98

ZW

PM and disk change out. The 1a SONIC would not respond to operator command’s. The 2b SONIC is working fine. The 1c Hygrometer would not respond to operator command’s. The 2d Hygrometer is also working fine at this time. Because of lack of spare C1 SONIC or Hygrometer no corrective action was taken at this time.

10/27/98

MK

The HYGROMETER is not responding to operator commands. Attempted to send an <esc> command to the HYGROMETER to reset it, no response. P/C HYGROMETER at 1940 GMT, no response again. Unable to replace due to lack of spares. RTS at 2000 GMT.

10/22/98

ZW

PM. The 60M ECOR is down at this time because of no spare parts.

10/15/98

MW/MR

There was no hygrometer installed at the 60M level. Lowered the boom at 1330 GMT. Installed *hygrometer on the 60M level. Was able to communicate with the sonic sensor and hygrometer. Raised the boom at 1505 GMT. Returned to service fully operational. *New - 940214

10/14/98

MW/MR

The instrument mentor wanted to know the status of this instrument. The C1 DCU is operational. The 60M ECOR sensor suite contains an operational sonic sensor but no hygrometer is installed. Did not drop the booms today due to high level winds. Will drop the booms first thing tomorrow morning and install a new hygrometer on the 60M level. Changed the 60M Bernoulli disk out at 1955 GMT since there was only 9.9MB of disk space left.

9/25/98

ZW

PM. The 60M ECOR is down at this time.

8/27/98

ZW

PM was not done at this time because the 60M ECOR is down at this time.

7/17/98

MR/KT

The 60M Hygrometer has not operated properly since installing it on 5/22. The sensor had not yet been removed due to lack of spares. Removed the *60M Hygrometer while the booms were lowered for T/H calibrations at 1445 GMT on 7/17/98. *Old - 940214

7/16/98

7/17/98

CAM/DC

All tower booms were lowered from 1458 GMT on 7/16/98 until 1513 GMT on 7/17/98 for 25M and 60M T/H calibrations. The booms were powered at ground level during that time.

7/1/98

ZW

PM and disk change out. The 1a SONIC and the 2d Hygrometer would not respond to operator commands. Power cycled the 1a SONIC and the 2d Hygrometer. the SONIC and Hygrometer would not respond to operator commands. Power cycled the ECOR system at 1945 GMT. The a1 SONIC and 2d Hygrometer still would not respond to operator commands form the laptop computer. Because of no spare sensors no more corrective action was taken. RTS 2000 GMT.

6/18/98

ZW

PM and disk change out. The 2b SONIC is working good, however the 2d Hygrometer would not respond to operator commands from the laptop computer. Power cycled several times with no success. Because of lack of spare Hygrometer no further corrective action was taken at this time. RTS 1945 GMT.

6/4/98

ZW

PM and disk change out. The 60M SONIC is working in good order, however the 60M Hygrometer would not respond to operator commands from the laptop computer. Power cycled the Hygrometer at 2000 GMT. Still could not communicate with the Hygrometer. Due to lack of spares no more corrective action taken at this time. RTS 2015 GMT.

5/22/98

ZW

On the last PM an existing problem with the 60M ECOR Hygrometer was found. Since the installation of the 60M Hygrometer the Q value was 0.00, and the H value was 01612 or in the same range. Because of a lack of spare no corrective action was taken. One Hygrometer was found in the other road crew's road supply. At 1400 GMT Site Ops lowered the 60M tower and R/R the 60M *Hygrometer. Attempted to monitor the new Hygrometer the laptop computer would give CE4. Power cycled the system several times but the CE4 did not change. The problem may be with the laptop computer program. The status will be checked on next PM. RTS 1500 GMT. * Old - 940212 *New - 940214

5/21/98

ZW

PM and disk change out. The 60M Hygrometer is still reporting bad data, however there are no spare Hygrometers at this time. The SONIC is still good. The DCU would not respond to operator commands from the Laptop computer. Power cycled the DCU at 1845 GMT. The communications with the DCU were restored. RTS 1900 GMT.

5/7/98

ZW

PM and disk change. The DCU would not respond to Operator commands from the laptop computer. Power cycled DCU at 1910 GMT. Communications restored. The 60M SONIC is working fine, however the Hygrometer readings are low Q is reading 0.0 and H is reading 01671. Because the spare Hygrometer stock is so low no replacement was made at this time. RTS 1925 GMT.

4/24/98

ZW

The 60M SONIC, and Hygrometer have been missing while the SONIC has been out for repairs. The 60M SONIC was returned from Applied Technologies. Powered off the DCU at 1415 GMT. Lowered the 60M instrument platform installed *SONIC and **Hygrometer. Raised the platform and powered on the DCU at 1504 GMT. The new SONIC is operating fine, however the new Hygrometer Q reading is 0.00. Temp, and pressure are good, but the H reading is 01630. There is a shortage of spare Hygrometers at this time so for now Hygrometer S/N 940212 which was also just out of the box from Applied Technologies will stay on the 60M platform operating and will be checked on next PM. RTS 1504 GMT. *New - 940405 **New - 940212

4/23/98

ZW

PM. The SONIC, and Hygrometer are still missing at this time.

4/10/98

ZW

PM the 60M SONIC, and Hygrometer are still missing because of no spare SONIC.

3/26/98

ZW

PM and disk change out. The 25M SONIC is the only sensor operational at this time. Due to lack of spare 60M SONIC which should be back any time from Applied Technologies.

3/13/98

ZW

On the PM if was found that the SONIC would not respond to operator commands from the laptop. After power cycling the SONIC was not successful site Ops removed the *SONIC at 1900 GMT. No spare C1 SONIC at this time. So the 60M part of the C1 ECOR is down at this time. The Hygrometer Q reading was very high, and the H was reading 01523 which is to low. After power cycling the Hygrometer was not successful site Ops removed the **Hygrometer at 1900 GMT. No spare C1 SONIC at this time which is necessary to operate the hygrometer, So the 60M part of the C1 ECOR is down at this time. RTS 2100 GMT. *Old - 940405 **Old - 940216

2/27/98

ZW

PM and disk change out. The 60M Sonic, and Hygrometer are working fine. RTS 1910 GMT.

2/13/98

ZW

The PM on the C1 ECOR was yesterday. It was discovered at this time that the DCU would not respond to operator commands. Attempts to power cycle the DCU would not work because the Bernoulli disk would not stop spinning. The 9 pin connector where the laptop connects to the DCU also needs to be replaced. At 1930 Site Ops powered off the C1 ECOR DCU. *R/R DCU with one just back from Applied Technologies INC. The new Hygrometer worked fine. RTS 2030 GMT * Old - 960201 * New - 940101

2/12/98

ZW

PM and disk change. The C1 ECOR DCU would not respond to operator commands. Powered the DCU off waited 10 sec. and powered it back on. The Bernoulli disk would not stop spinning. Site Ops attempted to power cycle the DCU several more times with no success, Because each time the disk would not stop spinning. Site Ops left the Bernoulli spinning, and returned about 1 hour later. The Bernoulli had stopped spinning but the DCU would not communicate. Power cycled the DCU again, and again the Bernoulli would not stop spinning. Site Ops left the system with the Bernoulli still spinning. RTS 1835 GMT.

1/29/98

ZW

PM and disk change. Status ok. RTS 2115 GMT.

1/16/98

SS

PM and disks were changed out. The 60M SONIC readings for U, V, W, and temp are 9999, and the status reading is 0000. Power cycled the System three time but the problem remains. RTS 1450 GMT.

12/31/97

ZW

PM and Disk change. Power cycled the ECOR at 1900 GMT. Status ok. RTS 1920 GMT.

12/12/97

ZW

PM. The C1 ECOR *DCU Bernoulli disk drives would not stop spinning. This is required to communicate with the DCU. R/R the disk drives which did not work. Removed the CPU board from the DCU which was operating the C1 ECOR at the time S/N 940101, and placed the CPU board in another double drive DCU S/N 960201 which came from the ECOR at E24 Cyril . The DCU From E24 Cyril S/N 960201 is working in the C1 ECOR at this time. The DCU from C1 ECOR will be sent back to Applied Technologies. The 60M ECOR Hygrometer was reporting Q 0.00, and H 01559 Which was not performing properly in a comparison to the 60M Hygrometer. Removed and Replaced The 60M ECOR **Hygrometer at 1800 GMT. The new Hygrometer is operating fine. Status ok. RTS 1830 GMT. *Old - 940101 *New - 960201 **Old - 960202 **New - 940216

12/5/97

ZW

PM. DCU still missing. System is still down.

11/5/97

CAM

After installing a new DCU on 10/24/97, the 25/60m ECOR could not be monitored. The Bernoulli disks spun continuously. Due to no spare DCU's and mentor being expected on site 11/4, the system was left as is for mentor to examine. Mentor arrived on 11/4 and the booms were lowered for maintenance. As expected, the system could not be monitored. Installed a video card in the DCU and hooked it up to a monitor. By monitoring the boot up sequence and diagnostics, it was determined that the DCU failed while trying to read drive D which is one of the Bernoulli drives. No spare Bernoulli drives on hand at this time. System remains off line until a spare is obtained.

10/24/97

ZW

PM. The C 1 ECOR was down because the DCU was sent back to Applied Technologies to be fixed. Installed *DCU. However the problem remains the same. The DCU will not communicate with the laptop computer. The Bernoulli disk for the 25 meter, and the Bernoulli disk for the 60 meter will not stop spinning. Communication with the DCU can not be established until the disks stop spinning. No corrective action taken at this time site ops will bring back the only other DCU with two Bernoulli disk drives for a possible alternate for the C 1 ECOR. *New - 940101

10/10/97

ZW

This instrument is not operational at this time.

9/12/97

ZW

PM and removed disks. The DCU would not respond to operator commands from the laptop computer. Power cycled the DCU several times to no avail. *Removed DCU. This instrument is down until a spare is available. * Old - 940101

8/15/97

ZW

PM and disks change out. The 25 meter SONIC and Hygrometer are missing. No corrective action taken due to lack of spare. The 60 meter SONIC and Hygrometer are working fine. RTS 1350 GMT.

8/14/97

GRS

Raised the 60 boom at 1602 GMT. RTS ECOR

8/13/97

GRS

Lowered the 60m boom at 1600 GMT to perform calibrations on the temp/hum sensors. The boom will remain down until the temp/hum calibrations are complete.

8/5/97

MW

Was not able to establish communications with the 60M Hygrometer using a laptop computer. Powered the DCU off at 1758 GMT. Lowered the 60M boom at 1803 GMT. *R/R Hygrometer. Raised the 60M boom at 2000 GMT. Powered the DCU on at 2010 GMT. RTS at 2015 GMT. * Old – 940210 * New - 960202

7/18/97

ZW

PM and disks were changed. The power cycled DCU at 1730 GMT. Communications restored. The 1A SONIC in not working at this time. No spare C1 SONIC at this time. Other than that Status is ok. RTS 1740 GMT.

7/3/97

ZW

PM and disks change. SONIC 1A is bad no spare at this time. Other than that status ok. RTS 1530 GMT.

6/20/97

ZW

PM and disk change. SONIC A1 would not respond to operator commands. Power cycled system at 1600 GMT with no positive results. Status other than that ok. RTS at 1615 GMT.

6/5/97

ZW

PM and disk change. Power cycled DCU at 1910 GMT. Power cycled SONIC 1A, 2B, and Hygrometer 1C, 2D at 1920. SONIC 1A is still bad however there are no spare at this time. Other than that status ok. RTS 1940 GMT.

5/23/97

ZW

PM and disks changed. SONIC 1A is bad however there are no spare at this time. Other than that Status ok. RTS 1330 GMT.

5/9/97

ZW/MK

PM and disk change out. Power cycled DCU at 1437 GMT. SONIC A was giving bad readings power cycled SONIC A at 1451 GMT to no avail. No spare at this time. Other than SONIC A Status is ok. RTS 1506 GMT.

5/01/97

CM/MW

Lowered boom at 1720 GMT. Installed repaired *DCU and **Sonic. Performed PM. Set DCU clock at 1835 GMT. Time delta was -00:04:17. Raised boom and RTS status O.K. at 1920 GMT. * New - 940101 ** New - 940405

2/18/97

MDW/CAM

Further T/S suggested DCU was bad. *Removed DCU and returned to vendor for repair. * Old - 940101

2/11/97

CM/MS

T/S found that the 60M Sonic power supply was bad, thus tripping the GFCI. * Removed Sonic sensor and electronics to return to vendor. Hygrometer left in place but will be inop since it also receives DC power from the Sonic electronics. Raised boom at 2120 GMT. * Old - 940405

2/10/97

MW/ZW

Lowered booms at 2000 GMT to install repaired * Sonic and ** Hygrometer. After sensors were installed the "B Upper" GFCI would not stay reset. Removed power and will T/S tomorrow. * New - 940405 ** New - 940210

10/11/96

MDW

Removed * Sonic and ** Hygrometer at 17:30 GMT due to power problems. Will send units back to Vendor for repairs. * Old - 940405 ** Old - 940216

9/30/96

MDW

PM and disk change. Could not communicate with the Hygrometer and Sonic. Power cycled the entire system at 1946 GMT. Changed Bernoulli from 9/13/96 to 9/30/96. No communications at 2000 GMT. Power cycled at 2000 GMT. RTS at 2015 GMT.

9/10/96

CAM/DKB

Lowered for 60m RH maintenance.

8/12/96

CAM/MDW

PM and data collection. Had to power cycle DCU and sensors while T/S 25M sonic. RTS at 1830 GMT.

7/8/96

CAM

PM & Data collection.

7/3/96

CAM/MW

Special visit. Lowered tower at 1925 GMT to T/S 25 ECOR Hygrometer. RTS at 2020 GMT.

6/24/96

MW/ZW

PM & Data collection. Added water to cool cell. RTS at 1815 GMT.

6/7/96

MW/ZW

PM & Data collection. RTS at 1800 GMT.